### v2

@@ 6.3 Analysis of the free time
Figure 12: The behavior of the the $\chi^{2}$ function and the KL divergence with time for $\gamma=0$ (top, left), $\gamma=0.3$ (top, right), $\gamma=0.5$ (bottom, left), and $\gamma=0.8$ (bottom, right). In each panel, the data points for $\chi^{2}$ lie above the data points for the KL divergence. We choose $s=1/2$, $L=8$, and take averages over $500$ realizations.
Figure 12 illustrates the temporal evolution of the $\chi^{2}$ function and the KL divergence for four distinct values of $\gamma$. The solid gray lines indicate the associated free times for each method, determined by fitting the following exponentially decaying function:
@@
Figure 13: (a) The free time (or the arcsine time for spin $1/2$) with the deformation $\gamma$, computed from the $\chi^{2}$ function (brown), KL divergence (red), and the Kolmogorov-Smirnov (KS) test (blue). (b) The finite-size scaling of free time for a fixed $\gamma=0$. The corresponding realizations are taken as $L=6\,(5000)$, $L=7\,(3000)$, $L=8\,(1000)$, $L=9\,(500)$, $L=10\,(200)$, $L=11\,(100)$, and $L=12\,(50)$. The dimension of the Hilbert space is $2^{L}$. The errors are computed by the standard error estimates in Mathematica.
@@
[…]
@@
[…]
The procedure of the KS test involves computing the statistic $x=\sqrt{N}D_{N}$ and evaluating whether it is significantly large. This is done by evaluating the associated $p$-value $=1-\text{CDF}_{\text{K}}(x)$, which is the probability of finding a larger value for the parameter $x=\sqrt{N}D_{N}$ given that the null hypothesis is true (the two distributions are the same). If the $p$-value is smaller than 0.05, one concludes that the parameter is too large and the null hypothesis is false, namely, the two distributions are not the same. Otherwise, if the $p$-value is larger than 0.05, one concludes that the null hypothesis is true and the two distributions are the same. [^17] Given two distributions, the KS test can be easily implemented in Mathematica by using the command KolmogorovSmirnovTest. Figure 15 illustrates the Kolmogorov PDF and the $p$-value associated with a particular value of $x=\sqrt{N}D_{N}$.
@@
We apply the KS test to assess whether the empirical distributions for the eigenvalues of $A(0)+A(t)$ follow the prediction from free probability. In Fig. 14, we show the $p$-value of the test as a function of time for different values of $\gamma$. The figure shows that at early times the $p$-value takes values smaller than 0.05, indicating that the empirical distribution and the free probability prediction are not the same. However, as time increases, the $p$-value increases, taking values larger than 0.05, indicating that the free probability prediction has emerged. The free time is computed as the time at which the $p$-value crosses the value 0.05.181818We also checked with other tests such as Pearson’s $\chi^{2}$ test and confirm the strikingly similar behavior of free time, as in Fig. 13.
@@ 7 Conclusion and outlook
Within the framework of free probability, we revisit the calculation of the DOS for the Rosenzweig–Porter (RP) random matrix ensemble and study the spectral statistics of operators, highlighting the approximate emergence of asymptotic freeness in this model. Since the RP model is defined as the sum of two independent random matrices, the two terms are free in the sense of free probability in the limit when the dimension of the matrices is sufficiently large. This enables the use of free probability techniques to compute the DOS. Using a general approach based on subordination functions, we derive an explicit expression for the DOS of the RP model, reproducing previous results obtained via the replica technique Venturelli:2022hka . Our analysis focuses on the fractal phase, where we express the DOS as perturbative corrections to the Gaussian profile characteristic of the localized phase. As illustrated in Fig. 3, the DOS evolves smoothly from a Wigner semicircle to a Gaussian distribution as the parameter $\gamma$ increases. In particular, during the transition from the ergodic to the fractal phase, the DOS resembles a perturbed Wigner semicircle. It would be interesting to extend the methodology developed here to compute the DOS in this intermediate regime.
It has been recently proposed that chaotic dynamics drives the components of a sum of two operators, one of which is time-evolved, toward asymptotic freeness with respect to non-evolved ones Camargo:2025zxr . However, the precise mechanisms underlying the emergence of freeness remain unclear. For instance, is it sufficient for the spectral statistics of the Hamiltonian to follow random matrix theory? What role do the eigenvectors play in this process? To address these questions, we studied the emergence of free probability predictions for the spectrum of operators of the form $S_{1}+S_{L}(t)$ in the RP model, an ensemble of random matrix Hamiltonians where the rotational invariance of the classical Gaussian ensembles is broken explicitly by introducing a parameter, and depending on the value of this parameter the energy eigenvectors show different types of localization properties. Our analysis reveals that the transition from classical convolution to free convolution manifests exclusively in the ergodic phase. In contrast, neither the fractal phase nor the localized phase exhibits this free convolution behavior.
@@
We have utilized various measures to accurately characterize the timescale at which free probability predictions emerge, marking the onset of free time. These methods include distance metrics like the $\chi^{2}$ function and KL divergence, alongside statistical techniques such as hypothesis testing. Notably, all methods yield consistent and reliable results. Additionally, the agreement between the KS test (based on the CDF) and both the KL divergence and $\chi^{2}$ test (based on the PDF) underscores the robustness of PDF-based methods in our analysis. This consistency confirms that the choice of histogram bin size does not affect our findings within the estimated error ranges, a conclusion supported by the considerable number of Hamiltonian ensembles in the operator statistics evaluation.
@@
In the present study, our focus has been restricted to the statistical properties of eigenvalues. However, the role of eigenvectors is equally crucial, especially in the context of thermalization. A natural extension would be to investigate the localization and delocalization behavior of the eigenvectors of the operator $S_{1}+S_{L}(t)$, which could shed light on the mechanisms underlying the ETH. Such an investigation would be especially illuminating when approached through the framework of free cumulants, potentially revealing interesting connections with recent developments in the field Pappalardi:2022aaz ; Pappalardi:2023nsj ; Fritzsch:2024qjn ; Wang:2025mzl ; Vallini:2024bwp ; Fava:2023pac , with our results on operator statistics. Especially, it would be an excellent direction to compare the free time studied here with the timescale for the onset of freeness introduced in Ref. Vallini:2024bwp , generalizing to systems hosting multifractal states Khaymovich:2021tkj . Pursuing this direction would require the consideration of finite-temperature states, thereby necessitating a generalization of the current statistical framework. Addressing these questions will be the subject of future investigations.
Acknowledgements
We thank Yichao Fu, Koji Hashimoto, Kohei Kawabata, Tokiro Numasawa, Silvia Pappalardi, Tanay Pathak, Hal Tasaki, and Masafumi Udagawa for fruitful discussions. We also thank Jonah Kudler-Flam for useful comments on the draft. P.N. acknowledges the hospitality of Gwangju Institute of Science and Technology (GIST), Gakushuin University and the Institute of Solid State Physics (ISSP), University of Tokyo, during the final stages of the work. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791, NRF-2020R1I1A2054376), the framework of international cooperation program managed by the NRF of Korea (RS-2023-NR119907, RS-2024-00440725), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the NRF funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. V.J. and H.C. are supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). The work of P.N. is supported by the JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05190.
@@ Appendix A Comparison of Haar random unitaries and exponential of Gaussian ensembles
Figure 16: Nearest-level spacing distribution for the unitaries $U_{\mathrm{COE}}$ (left), $U_{\mathrm{CUE}}$ (middle) and $U_{\mathrm{GXE}}$ (right), with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ (both produce identical plots) for $s=1/2$. The red lines are the analytic results for $\upbeta=1,2$ and 00 respectively (from left to right). We choose $L=8$ and average over $5000$ samples. For $U_{\mathrm{GXE}}$, we choose $t=50$.
Figure 17: Histogram of the $r$-ratio for the unitaries $U_{\mathrm{COE}}$ (left), $U_{\mathrm{CUE}}$ (middle) and $U_{\mathrm{GXE}}$ (right), with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ (both produce identical plots) for $s=1/2$. The red lines are the analytic results for $\upbeta=1,2$ and 00 respectively (from left to right). We choose $L=8$ and average over $5000$ samples. For $U_{\mathrm{GXE}}$, we choose $t=50$.
@@
Appendix B Free additive convolution of spin-$s$ operators

### v3

@@ 6.3 Analysis of the free time
Figure 14: The behavior of the $\chi^{2}$ function and the KL divergence with time for $\gamma=0$ (top, left), $\gamma=0.3$ (top, right), $\gamma=0.5$ (bottom, left), and $\gamma=0.8$ (bottom, right). In each panel, the data points for $\chi^{2}$ lie above the data points for the KL divergence. We choose $s=1/2$, $L=8$, and take averages over $500$ realizations.
Figure 14 shows the temporal evolution of the $\chi^{2}$ function and the KL divergence for four representative values of $\gamma$. The solid gray lines mark the corresponding free times, obtained in each case by fitting the data to the following exponentially decaying function:
@@
Figure 15: (a) The free time (or the arcsine time for spin $1/2$) with the deformation $\gamma$, computed from the $\chi^{2}$ function (brown), KL divergence (red), and the Kolmogorov-Smirnov (KS) test (blue). (b) The finite-size scaling of free time for a fixed $\gamma=0$. The corresponding realizations are taken as L=6(5000)L=6\,(5000), L=7(3000)L=7\,(3000), L=8(1000)L=8\,(1000), L=9(500)L=9\,(500), L=10(200)L=10\,(200), L=11(100)L=11\,(100), and L=12(50)L=12\,(50). The dimension of the Hilbert space is $2^{L}$. The errors are computed by the standard error estimates in Mathematica.
@@
$\text{CDF}_{\text{ref}}(\lambda)=\int_{-\infty}^{\lambda}\text{PDF}_{\text{ref}}(\lambda^{\prime})\,d\lambda^{\prime}\,.$ (72)
@@
$\text{CDF}_{\mathrm{K}}(x)=\int_{-\infty}^{x}\text{PDF}_{\mathrm{K}}(x^{\prime})\,dx^{\prime}=1-2\sum_{n=1}^{\infty}(-1)^{n-1}e^{-2n^{2}x^{2}}\,$, (77)
The procedure of the KS test involves computing the statistic $x=\sqrt{N}D_{N}$ and evaluating whether it is significantly large. This is done by evaluating the associated $p$-value $1-\text{CDF}_{\text{K}}(x)$, which is the probability of finding a larger value for the parameter $x=\sqrt{N}D_{N}$ given that the null hypothesis is true (the two distributions are the same). If the $p$-value is smaller than 0.05, one concludes that the parameter is too large and the null hypothesis is false, namely, the two distributions are not the same. Otherwise, if the $p$-value is larger than 0.05, one concludes that the null hypothesis is true and the two distributions are the same. [^18] Given two distributions, the KS test can be easily implemented in Mathematica by using the command KolmogorovSmirnovTest. Figure 17 illustrates the Kolmogorov PDF and the $p$-value associated with a particular value of $x=\sqrt{N}D_{N}$.
@@
We apply the KS test to assess whether the empirical distributions for the eigenvalues of $A(0)+A(t)$ follow the prediction from free probability. In Fig. 16, we show the $p$-value of the test as a function of time for different values of $\gamma$. The figure shows that at early times the $p$-value takes values smaller than 0.05, indicating that the empirical distribution and the free probability prediction are not the same. However, as time increases, the $p$-value increases, taking values larger than 0.05, indicating that the free probability prediction has emerged. The free time is computed as the time at which the $p$-value crosses the value 0.05.191919We also checked with other tests such as Pearson’s $\chi^{2}$ test and confirm the strikingly similar behavior of free time, as in Fig. 15.
@@ 7 Conclusion and outlook
Within the framework of free probability theory, we have revisited the calculation of the DOS for the Rosenzweig–Porter (RP) random matrix ensemble and studied the spectral statistics of operators to understand the emergence of asymptotic freeness in different phases of this model. Since the RP model is defined as the sum of two independent matrices, one of which, in the standard version of this model, is drawn from the classical Gaussian random matrix ensembles, the two terms are free according to the free probability theory in the limit when the dimension of the matrices is sufficiently large. This enables us to use free probabilistic techniques to compute the DOS of this model. Using a general approach based on subordination functions, we derive an explicit expression for the DOS of the RP model in the fractal phase, reproducing previous results obtained via the replica technique Venturelli:2022hka . Our analysis focuses on the fractal phase, where we could write the DOS as perturbative corrections to the Gaussian profile characteristic of the localized phase. As illustrated in Fig. 3, the DOS evolves smoothly from a Wigner semicircle to a Gaussian distribution as the parameter $\gamma$ increases. In particular, during the transition from the ergodic to the fractal phase, the DOS resembles a perturbed Wigner semicircle. It would be interesting to extend the methodology developed here to compute the DOS in this intermediate regime.
It has been recently proposed that chaotic dynamics drives the components of a sum of two operators, one of which is time-evolved, toward asymptotic freeness with respect to non-evolved ones Camargo:2025zxr ; Chen:2024zfj . However, the precise mechanisms underlying the emergence of freeness remain unclear. For instance, is it sufficient for the spectral statistics of the Hamiltonian to follow RMT? What role do the eigenvectors play in this process? To address these questions, we studied the emergence of free probability predictions for the spectrum of operators of the form $S_{1}+S_{L}(t)$ in the RP model, an ensemble of random matrix Hamiltonians where the rotational invariance of the classical Gaussian ensembles is broken explicitly by introducing a parameter, and depending on the value of this parameter the energy eigenvectors show different types of localization properties. Our analysis reveals that the transition from classical convolution to free convolution manifests exclusively in the ergodic phase, where the eigenvectors are fully delocalized. In contrast, for neither the fractal phase nor the localized phase do the operator spectral statistics match the free convolution prediction.
@@
To substantiate the above conclusions, in Appendix B we study the eigenvalue statistics of the same time-evolved operators where the time evolution is now generated by a Hamiltonian drawn from a decorrelated version of the RP ensemble. The characteristic feature of the decorrelated ensembles (also known as the Poissonian ensembles in the literature Magan:2024nkr ) constructed from a given physical ensemble (or a quantum many-body system) is that a Hamiltonian drawn from it has the same eigenvector localization/delocalization/fractal property as that of the original ensemble, while, if there are correlations between the eigenvalues in the original ensemble they are removed by hand when constructing the decorrelated ensemble. Our results, presented in Appendix B, show that the decorrelated version of the RP ensemble gives rise to the same eigenvalue statistics of the time-evolved operators as those of the original RP model, once again indicating that complete delocalization of eigenvectors is sufficient for producing eigenvalue statistics compatible with free probability, not eigenvalue correlation.
In section 6.3, we have utilized various measures to accurately characterize the timescale at which free probability predictions emerge, marking the onset of free time. These methods include distance metrics like the $\chi^{2}$ function and KL divergence, alongside statistical techniques such as hypothesis testing. Notably, all methods yield consistent and reliable results. Additionally, the agreement between the KS test (based on the CDF) and both the KL divergence and $\chi^{2}$ test (based on the PDF) underscores the robustness of PDF-based methods in our analysis. This consistency confirms that the choice of histogram bin size does not affect our findings within the estimated error ranges, a conclusion supported by the considerable number of Hamiltonian ensembles in the operator statistics evaluation.
@@
In the present study, our focus has been restricted to the statistical properties of eigenvalues of an operator of the form $S_{1}+S_{L}(t)$. However, the role of eigenvectors is equally crucial, especially in the context of thermalization. A natural extension would be to investigate the localization and delocalization behavior of the eigenvectors of the operator $S_{1}+S_{L}(t)$, which could shed light on the mechanisms underlying the ETH. Such an investigation would be especially illuminating when approached through the framework of free cumulants, potentially revealing interesting connections with recent developments in the field Pappalardi:2022aaz ; Pappalardi:2023nsj ; Fritzsch:2024qjn ; Wang:2025mzl ; Vallini:2024bwp ; Fava:2023pac , with our results on operator statistics. Especially, it would be an excellent direction to compare the free time studied here with the timescale for the onset of freeness introduced in Ref. Vallini:2024bwp , generalizing to systems hosting multifractal states Khaymovich:2021tkj . Pursuing this direction would require the consideration of finite-temperature states, thereby necessitating a generalization of the current statistical framework. Addressing these questions will be the subject of future investigations.
8 Acknowledgements
We thank Neil Dowling, Yichao Fu, Koji Hashimoto, Kohei Kawabata, Ivan M. Khaymovich, Tokiro Numasawa, Silvia Pappalardi, Tanay Pathak, Hal Tasaki, and Masafumi Udagawa for fruitful discussions. We also thank Jonah Kudler-Flam for useful comments on the draft. P.N. acknowledges the hospitality of Gwangju Institute of Science and Technology (GIST), Gakushuin University, and the Institute of Solid State Physics (ISSP), University of Tokyo, during the final stages of the work. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791), the framework of international cooperation program managed by the NRF of Korea (RS-2025-02307394), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. RS-2023-NR068116) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. This research was also supported by the Regional Innovation System & Education (RISE) program through the (Gwangju RISE Center), funded by the Ministry of Education (MOE) and the (Gwangju Metropolitan City), Republic of Korea (2025-RISE-05-001). V.J. and H.C. are supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). The work of P.N. is supported by the JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05190.
@@ Appendix A Comparison of Haar random unitaries and exponential of Gaussian ensembles
Figure 18: Nearest-level spacing distribution for the unitaries $U_{\mathrm{COE}}$ (left), $U_{\mathrm{CUE}}$ (middle) and $U_{\mathrm{GXE}}$ (right), with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ (both produce identical plots) for $s=1/2$. The red lines are the analytic results for $\upbeta=1,2$ and 0 respectively (from left to right). We choose $L=8$ and average over $5000$ samples. For $U_{\mathrm{GXE}}$, we choose $t=50$.
Figure 19: Histogram of the $r$-ratio for the unitaries $U_{\mathrm{COE}}$ (left), $U_{\mathrm{CUE}}$ (middle) and $U_{\mathrm{GXE}}$ (right), with $\mathrm{X}\in\{\mathrm{O},\mathrm{U}\}$ (both produce identical plots) for $s=1/2$. The red lines are the analytic results for $\upbeta=1,2$ and 0 respectively (from left to right). We choose $L=8$ and average over $5000$ samples. For $U_{\mathrm{GXE}}$, we choose $t=50$.
@@
Appendix B Role of eigenvectors in freeness and decorrelated ensembles
As discussed in Sec. 6, the eigenvalue statistics of operators of the form $S_{1}+S_{L}(t)$, with time evolution governed by a Hamiltonian drawn from the RP ensemble, coincide with the predictions of free probability only in the ergodic phase, where eigenvectors are fully delocalized. In contrast, in the fractal phase, although eigenvalue correlations persist, the corresponding statistics do not approach the free-probability limit even at late times, reflecting the fractal character of the eigenstates. To isolate the respective contributions of eigenvector delocalization and eigenvalue correlations in the emergence of freeness, we consider a variant of App. A. Specifically, we analyze operator statistics under time evolution generated by Hamiltonians drawn from a decorrelated version of the RP ensemble, building on previous work Camargo:2025zxr where such a situation was considered for classical Gaussian random matrix ensembles.
In essence, the decorrelated ensemble is constructed from a given random-matrix ensemble such that a Hamiltonian $H$ drawn from it preserves the localized, delocalized, or fractal structure of the original eigenvectors, while eliminating any correlations among the eigenvalues Nandy:2024zcd ; Magan:2024nkr . Concretely, consider a realization of matrices $A$ and $B$ sampled according to Eq. (20), yielding an RP Hamiltonian $H$ with eigenvectors $\lvert n\rangle$ and eigenvalues $E_{n}$ (ordered in ascending $n$). The spectrum defines a discrete density of states (DOS) $\rho(E)$, whose ensemble average $\overline{\rho(E)}$ approaches the $\gamma$-dependent DOS in the thermodynamic limit $N\to\infty$ (see Fig. 3). While the true eigenvalues $E_{n}$ fluctuate around $\overline{\rho(E)}$ and exhibit $\gamma$-dependent correlations, one may instead draw $N$ independent samples $E_{n}^{\prime}$} from $\overline{\rho(E)}$. These “decorrelated” eigenvalues reproduce the same average DOS but, by construction, lack the original correlations and follow Poissonian level statistics for any $\gamma$ Nandy:2024zcd . A new Hamiltonian $H^{\prime}$ can then be defined by combining the original eigenvectors $\lvert n\rangle$ with the decorrelated eigenvalues $E_{n}^{\prime}$, ordered ascendingly. This procedure can be repeated for arbitrary realizations of the RP ensemble, since every such Hamiltonian is diagonalizable with real eigenvalues.
Figures 21-23 show the operator spectral statistics of the spin-$1/2$ operator sum $S_{1}+S_{L}(t)$ at different times for values of $\gamma$ corresponding to the ergodic ($\gamma=0$), fractal ($\gamma=1.5$), and localized ($\gamma=5$) phases, respectively. In the ergodic phase, the continuous DOS is the Wigner semicircle. For the fractal and localized regimes, we employ the corrected DOS obtained in Sec. 4.4, up to $O(\alpha^{2})$ with $\alpha=N^{-\gamma/2}$, as given in Eq. (4.3). This choice is appropriate since the values of $\gamma$ considered satisfy $\gamma>1$, consistent with the perturbative scheme discussed in Sec. 4.3.
Refer to caption Refer to caption
Figure 21: Statistics of eigenvalues of the sum of spin-$1/2$ operators $S_{1}(0)+S_{L}(t)$ for increasing values of time under time evolution by a Hamiltonian drawn from the RP ensemble in the ergodic phase $\gamma=0$ (yellow) and under time evolution by a Hamiltonian drawn from a decorrelated ensemble (gray). The overlap between these two distributions is shown in dark brown. The blue curve is the arcsine distribution. The bin size in each histogram is $\{0.06\}$, $L=8$, and the number of independent realizations considered is $500$.
Refer to caption Refer to caption
Figure 22: Statistics of eigenvalues of the sum of spin-$1/2$ operators $S_{1}(0)+S_{L}(t)$ for increasing values of time under time evolution by a Hamiltonian drawn from the RP ensemble in the fractal phase $\gamma=1.5$ (yellow) and under time evolution by a Hamiltonian drawn from a decorrelated ensemble (gray). The overlap between these two distributions is shown in dark brown. The blue curve is the arcsine distribution. The bin size in each histogram is $\{0.06\}$, $L=8$, and the number of independent realizations considered is $500$.
Refer to caption Refer to caption
Figure 23: Statistics of eigenvalues of the sum of spin-$1/2$ operators $S_{1}(0)+S_{L}(t)$ for increasing values of time under time evolution by a Hamiltonian drawn from the RP class in the localized phase $\gamma=5$ (yellow) and under time evolution by a Hamiltonian drawn from a decorrelated ensemble (gray). The overlap between these two distributions is shown in dark brown. The bin size in each histogram is $\{0.06\}$, $L=8$, and the number of independent realizations considered is $500$.
We observe that the operator spectral statistics generated by time evolution under decorrelated ensembles closely track those obtained from the full RP Hamiltonian, consistent with the results of Ref. Camargo:2025zxr for Hamiltonians drawn from Hermitian Gaussian ensembles. Notably, this agreement holds even at intermediate timescales, before the statistics converge to their asymptotic distributions at late times. These results reinforce the conclusion of Ref. Camargo:2025zxr that the eigenvectors play the central role in determining the operator statistics of sums of the form $X_{i}(0)+X_{j}(t)$, and that in the ergodic case they suffice to reproduce the free-probability prediction regardless of eigenvalue correlations. In the fractal phase, however, we observe small deviations at intermediate times, whereas in the localized phase, the decorrelated and true statistics remain nearly indistinguishable across all timescales.
Appendix C Free additive convolution of spin-$s$ operators